OpenStudy (anonymous):

Use the elimination method to solve the system. 5x+3y=-6 3x-2y=4

OpenStudy (anonymous):

PLZ help will give medal

OpenStudy (texaschic101):

5x + 3y = -6 -->(2) 3x - 2y = 4 -->(3) ---------------- 10x + 6y = -12 (result of multiplying by 2) 9x - 6y = 12 (result of multiplying by 3) ----------------add x = 0 now sub 0 in for x in either of the original equations 5x + 3y = -6 5(0) + 3y = -6 0 + 3y = -6 3y = -6 y = -2 check... 3x - 2y = 4 3(0) - 2(-2) = 4 0 + 4 = 4 4 = 4(correct) x = 0 and y = -2 or (0,-2)

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

nope

OpenStudy (texaschic101):

good deal....glad I could help :)